OpenStudy (anonymous):

The linear approximation for the function f(x)=x^5 near x=2 is 32+80(x-2) Check this. A.) What number c will give the best quadratic approximation x^5 approximate to 32+80(x-2)+c(x-2)^2. With the help of myininaya I got c=80. B.)If this approximation is used for various x's in the interval [2,2.1], can you be certian that the error is no bigger than .05? Explain using Taylor's inequality (the Error Bound). C.) Graph part equation Part A now that we have c. On interval [2,2.1]. note: I don't need someone to graph this, I just want to make sure I have right approach. Would i just pick values like 2, 2.05, and 2.1 for x, set equation equal to y and solve to get points?

myininaya (myininaya):

f(x)=x^5 f'(x)=5x^4 f''(x)=___?__ f(2)=2^5=32 f'(2)=5(2)^4=5(16)=80 f''(2)=___?__ f(x) approx=f(a)+f'(a)(x-a) f(x) approx=f(a)+f'(a)(x-a)+f''(a)/2 * (x-a)^2 What is f''(a)/2?

OpenStudy (anonymous):

f"(x)= 20x^3 f"(2)=160 f"(a)=80 I'm not sure about the last part after f"(2)

myininaya (myininaya):

That is right. Now f''(a)/2=?

myininaya (myininaya):

wait f''(a)=160 not 80?

myininaya (myininaya):

You mean f''(a)/2=80 correct?

OpenStudy (anonymous):

yes